OpenStudy (anonymous):

can someone help me with question i don't know how to do it please and thank you! 2(n - 3) + 5 = 3(n - 1)

OpenStudy (anonymous):

2n-6=3n-3 3n-2n=6-3 n=3

OpenStudy (jamesj):

Step 1: Expand each side and simplify LHS (left hand side) = = 2(n-3) + 5 = 2n - 6 + 5 = 2n - 1 RHS = 3(n-1) = 3n - 3 Step 2: Set LHS = RHS and collect all n terms on one side 2n - 1 = 3n - 3 2n - 3n - 1 = 3n - 3n - 3 -n - 1 = -3 -n - 1 + 1 = -3 + 1 -n = -2 Step 3: Solve explicitly for n -n = -2 hence n = 2

OpenStudy (karatechopper):

rob and jesus r correct

OpenStudy (anonymous):

distribute on both sides.. 2n - 6 + 5 = 3n -3 add like terms... 2n -1 = 3n -3 bring vairables all to one side. 2n - 3n = 1 - 3 -n = -2 multiply both sides by -1 n = 2!

OpenStudy (jamesj):

Step 4: Verify result by substituting back into equation LHS = 2(n - 3) + 5 = 2(2 -3) + 5 = -2 + 5 = 3 RHS = 3 (n - 1) = 3(2 -1) = 3(1) = 3 Hence LHS = RHS and therefore n = 2 is indeed the solution.

OpenStudy (jamesj):

And for the record n = 3 is NOT a solution.

OpenStudy (anonymous):

Silly number 5! James is correct! I lost my 5 in the equation. n=2

OpenStudy (anonymous):

2(n - 3) + 5 = 3(n - 1) 2n-6+5=3n-3 2n-1=3n-3 2n-3n=-3+1 -n=-2 n=2